# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 2. We show that
2. Now, by item 1 of Proposition 3.1 and by Proposition 2.31, we have
$A_{h}(x)=C_{h}(x)$
and argue as in the previous case for $\xi=1$. $\Box$
This theorem generalizes the case $d\geq 3$ of Pólya’s theorem with $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.7.
Theorem 3.6 (5: $v=1$, con, $U(1)_{2}$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $B_{h}(1)_{2}\neq 0$ and
@@
2. In the previous case we have for $\xi=1$ that $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. The stated formula follows. $\Box$
Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=\infty$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1\,.$
2. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a 0-convex light weight, $B_{h},C_{h}\in H_{1}$, and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}F_{(A_{h})_{1/\xi}}(x)=1\,.$
Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have
$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}\,.$}
By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
By Proposition 2.31 and the subseries bound principle,
$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}c_{n}^{h}\leq\xi^{-n}d_{n}^{h}=1$
for every $n\in\mathbb{N}_{0}$. Hence $B_{h},C_{h}\in H_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
$(C_{h})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1\,,}$
where in the first and third equality we used Theorem 2.18. The stated formula follows.
2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
$\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,,}$
where in the second equality we used Theorem 2.17. The stated formula follows at once. $\Box$
4 Extensions of Pólya’s theorem with $v\neq 1$
In Theorems 4.34.10 we establish the second group of generalizations of Pólya’s theorem, in the more complicated case when the vertex to be visited by walks differs from their starting vertex. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, which has so far been irrelevant. For $\overline{v}\neq\overline{0}$, vertex-transitivity becomes relevant. We generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$. We are interested in the generating function
$A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
In the previous section we introduced the generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. Now we define
$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b}(v,\,n))x^{n}\,,}$
$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c}(v,\,n))x^{n}\,\text{ and}}$
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e}(v,\,n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. The constant coefficients are $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
We obtain an analog of Proposition 3.1. Since the proof is similar, we omit details and argue only on the level of semiformal method.
@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be the above generating functions. The following relations hold.
@@
    $C_{h,v}(x)=B_{h,v}(x)E_{h,v}(x)$.
Proof. 1. Walks $w$ weight-counted by $A_{h,v}(x)$ split at the first visit of $v$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
3. We consider walks $w$ weight-counted by $C_{h,v}(x)$. They start at $1$, end at $v$, and between avoid $v$. These walks uniquely split at the last visit of $1$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$. $\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then $a_{n}^{h,v}=\sum_{w\in W(v,n)}h(w)=\sum_{w\in W(v,n)}0=0$ for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight. Henceforth, we therefore assume that $v\in V(h)$. For $w,z\in\mathbb{C}$ we define the set
@@
Before we turn to generalizations of Pólya’s theorem in the case $v\neq 1$, we utilize the square in item 2 of Proposition 4.1. If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then in general the sum $\sum_{x\in Y}h(x)$ of the subseries need not be smaller than the sum $\sum_{x\in X}h(x)$. However, due to the mentioned square, we obtain the next monotonicity result for the sums
$b_{n}^{h,v}=\sum_{w\in W_{b}(v,\,n)}h(w)\,\text{ and }\,b_{n}^{h}=\sum_{w\in W_{b}(n)}h(w)\,.$}
It is expressed in terms of the generating functions $B_{h,v}(x)=\sum_{n\geq 0}b_{n}^{h,v}x^{n}$ and $B_{h}(x)=\sum_{n\geq 0}b_{n}^{h}x^{n}$.
@@
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{h,v},B_{h},C_{h,v}\in H_{1}$. Then for every real $c\in(-1,1)$ the following holds.
@@
    If $B_{h}(c)\neq 0$ then
    $\frac{B_{h,v}(c)}{B_{h}(c)}\leq 1\,.$
Proof. This follows from relations 2 in Proposition 4.1. $\Box$
We proceed to the remaining generalizations of Pólya’s theorem.
Theorem 4.3 (7: $v\neq 1$, gen, $U(1)_{1}$)
If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By Proposition 2.20, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. Thus by Proposition 2.22, $\big(C_{h,v}(1)_{1}\big)^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
If we replace the assumption of non-vanishing of $B_{h}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
Theorem 4.4 (7: $v\neq 1$, gen, $U(1)_{1}$)
If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
Proof. If $B_{h}(1)_{1}\neq 0$, we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1, and Propositions 2.20 and 2.19 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.20 give $C_{h,v}(1)_{1}=0$. Finally, item 1 of Proposition 4.1 and Proposition 2.20 give $A_{h,v}(1)_{1}=0$. $\Box$
Theorem 4.5 (8: $v\neq 1$, gen, $U(1)_{2}$)
Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $B_{h,v}(1)_{2}$ and $D_{h}(1)_{2}$ exist. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}

### v8

@@ 3.3 Pólya’s theorems for $v=1$ and $\xi$-convex weights
2. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. We use the displayed formula for $(C_{h})_{1/\xi}(1)_{1}$ with $\xi=1$. $\Box$
This theorem generalizes the case $d\geq 3$ of the formula (1) for $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.8.
Theorem 3.7 ($\exists\,U(1)_{2}$)
1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $\xi$-convex light weight. If $\sum_{n\geq 1}\big|h(B_{n})\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
2. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a 0-convex light weight. If $\sum_{n\geq 1}|h(B_{n})|<1$, then $B_{h}(1)_{2}\neq 0$ and
@@
2. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. We use the displayed formula for $(C_{h})_{1/\xi}(1)_{2}$ with $\xi=1$. $\Box$
Theorem 3.8 ($U(1)_{1}=\infty$)
1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $\xi$-convex light weight. If $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1\,.$
2. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ be a 0-convex light weight. If $B_{h}(x),C_{h}(x)\in H_{1}$ and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}(A_{h})_{1/\xi}(x)=1\,.$
Proof. 1. We proved this in Theorem 3.5.
2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Theorem 2.17,
$\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)=(B_{h})_{1/\xi}(1)_{1}=\pm\infty\,.$
It follows from this and the assumptions that
$\text{$(C_{h})_{1/\xi}(x)=1-\frac{1}{(B_{h})_{1/\xi}(x)}$ for every $x\in(1-\delta,\,1)$}\,.$}
We compute that
$\lim_{x\nearrow 1}(C_{h})_{1/\xi}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,.}$
In the second equality we used Theorem 2.17. The stated formula follows. $\Box$
4 General Pólya’s theorems for $v\neq 1$
In this last section, we generalize the results in Section 1.3 to weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We first introduce another three auxiliary generating functions $B_{v,h}(x)$, $C_{v,h}(x)$, and $E_{v,h}(x)$. The second identity in Proposition 4.2 contains a square, and in Theorem 4.4 we deduce a result on monotonicity from this. In Sections 4.2 and 4.3, we obtain general Pólya’s theorems for $v\neq 1$ with general and $\xi$-convex weights, respectively.
4.1 Auxiliary generating functions
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. As we know from Section 1.4, we are interested in the main generating function
$A_{v,\,h}(x)=\sum_{n\geq 0}h(A_{v,\,n})x^{n}\ \ (\in\mathbb{C}[[x]])$}
for the set $A_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$; $h(A_{v,n})$ is the sum of the series
$\sum_{w\in A_{v,\,n}}h(w)=\sum_{\begin{subarray}{c}w\in A_{v}\\ s(w)=n\end{subarray}}h(w)\,.$}
Besides the sets $A_{v}$, $B$, and $D$ of walks in $K_{\mathbb{N}}$, we use the set $B_{v}$ of walks in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$, the set $C_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and with the inner vertices different from $v$, and the set $E_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and with the inner vertices different from both $1$ and $v$. We define three additional auxiliary generating functions
$B_{v,\,h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}h(B_{v,\,n})x^{n},\ C_{v,\,h}(x)=\sum_{n\geq 0}h(C_{v,\,n})x^{n}\,\text{ and}}$
$E_{v,\,h(x)}$ $=$ ${\textstyle\sum_{n\geq 0}h(E_{v,\,n})x^{n}\,.}$
We have $h(A_{v,0})=h(C_{v,0})=h(E_{v,0})=0$ and $h(B_{0})=h(B_{v,0})=h(D_{0})=1$.
In the case $\overline{v}=\overline{0}$ the symmetry of the grid graph $G_{d}=\langle\mathbb{Z}^{d},E_{d}\rangle$, namely its vertex-transitivity, played no role, but in the case $\overline{v}\neq\overline{0}$ this symmetry is crucial; it ensures that the enumeration of walks does not depend on the starting vertex. The same holds for the general weighted case, and we introduce the weighted version of vertex transitivity.
Definition 4.1 ($v$-transitivity)
Let $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We say that the weight $h$ is $v$-transitive if there exists a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that
$f(1)=v\,\text{ and }\,\forall e\in\mathbb{N}_{2}:\;h(e)=h(f[e])\,.$
We obtain an analog of Proposition 3.1. The proofs are similar, and we only sketch them.
@@ 4.1 Auxiliary generating functions
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight, and $A_{v,h}(x)$, $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, $D_{h}(x)$, and $E_{v,h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
@@
    $C_{v,h}(x)=B_{v,h}(x)E_{v,h}(x)$.
Proof. 1. We split every walk $w\in A_{v}$ at the first visit of $v$ as $w=w_{1}w_{2}$. This splitting yields, when the $v$-transitivity of $h$ is applied to $w_{2}$, the stated identity.
2. It suffices to prove the first identity. We have the partition
$B=B_{v}\cup B_{v}^{\prime}$
where the set $B_{v}^{\prime}$ consists of the walks in $K_{\mathbb{N}}$ that start and end at $1$ and visit $v$. We split every walk $w\in B_{v}^{\prime}$ at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. The above partition, the reversal of walks (applied to $w_{3}$), and the $v$-transitivity of $h$ (applied to $w_{2}$) yield the stated identity.
3. We split every walk $w\in C_{v}$ at the last visit of $1$ as $w=w_{1}w_{2}$. This splitting yields, without help of the $v$-transitivity of $h$, the stated identity. $\Box$
One would expect to see for $v\neq 1$ the generating function of the set $F_{v}$ of walks in $K_{\mathbb{N}}$ going from $1$ to $v$. We managed for a long time without it, but at last we include two identities involving $F_{v}$. So we define another auxiliary generating function
$F_{v,\,h}(x)=\sum_{n\geq 0}h(F_{v,\,n})x^{n}\,.$}
Proposition 4.3
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight, and $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, and $F_{v,h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
  1. 1.
    $F_{v,h}(x)=C_{v,h}(x)B_{h}(x)$.
  2. 2.
    $B_{v,h}(x)=B_{h}(x)-\frac{F_{v,h}(x)^{2}}{B_{h}(x)}$.
Proof. 1. We split every walk $w\in F_{v}$ at the first visit of $v$ as $w=w_{1}w_{2}$. This splitting yields, when the $v$-transitivity of $h$ is applied to $w_{2}$, the identity.
$F_{v,\,h}(x)=C_{v,\,h}(x)+C_{v,\,h}(x)(B_{h}(x)-1)=C_{v,\,h}(x)B_{h}(x)\,.$
2. Using the reversal of walks and the $v$-transitivity of $h$, we get
$B_{h}(x)-\frac{F_{v,\,h}(x)^{2}}{B_{h}(x)}=B_{h}(x)-F_{v,\,h}(x)^{2}\sum_{j\geq 1}(-1)^{j-1}(B_{h}(x)-1)^{j-1}=B_{v,\,h}(x)\,.$}
Indeed, if $w$ is a walk in $K_{\mathbb{N}}$ from $1$ to $1$ with $k\in\mathbb{N}_{0}$ visits of the vertex $v$, then $w$ is counted exactly
$1-\sum_{j=1}^{k}\binom{k}{j}(-1)^{j-1}=\sum_{j=0}^{k}\binom{k}{j}(-1)^{j}=(1-1)^{k}\ \ (\in\{0,\,1\})$}
times in the left-hand side $\cdots$ of the displayed equality $\dots=B_{v,h}(x)$, which agrees with the count of $w$ in the right-hand side $B_{v,h}(x)$. $\Box$
Item 2 in the previous proposition and item 1 in this one provide two expressions for $C_{v,h}(x)$:
$C_{v,h}(x)^{2}=1-\frac{B_{v,\,h}(x)}{B_{h}(x)}\,\text{ and }\,C_{v,h}(x)=\frac{F_{v,\,h}(x)}{B_{h}(x)}\,.$}
Although the latter expression is simpler than the former, both in appearance and factually ($F_{v}$ has simpler definition than $B_{v}$), we use the former: due to inclusions $B_{v,n}\subset B_{n}$ we easily show the membership in $(0,1)$ in Theorem 4.9. In the simple latter expression we have disjoint sets $F_{v,n}$ and $B_{n}$, and it is not clear how to deduce that $(F_{v,h})_{1/\xi}(1)_{1}<(B_{h})_{1/\xi}(1)_{1}$.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then $h(w)=0$ for every walk $w\in A_{v}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight. Henceforth, we therefore always assume that $v\in V(h)$. For $w,z\in\mathbb{C}$, we define the set
@@ 2. Using the reversal of walks and the $v$-transitivity of $h$, we get
Before we proceed to general Pólya’s theorems for $v\neq 1$, we utilize the square in item 2 of Proposition 4.2. If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then the sum $h(Y)$ of the subseries need not be smaller than the sum $h(X)$. However, in the case of inclusions $B_{v,n}\subset B_{n}$, sums $h(B_{v,n})$ are, in a sense, indeed smaller than sums $h(B_{n})$, as shown in the following theorem.
@@
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{v,h}(x),B_{h}(x),C_{v,h}(x)\in H_{1}$. Then for every number $x\in(-1,1)$ the following holds.
@@
    If $B_{h}(x)\neq 0$ then
    $\frac{B_{v,\,h}(x)}{B_{h}(x)}\leq 1\,.$
Proof. By item 2 of Proposition 4.2, the assumptions, and Propositions 2.2, 2.4 and 2.6 we have for every number $x\in(-1,1)$ that
$B_{h}(x)=B_{v,h}(x)+C_{v,h}(x)^{2}B_{h}(x)\,.$
The claim 1 is clear. If $B_{h}(x)\neq 0$ then
$\frac{B_{v,\,h}(x)}{B_{h}(x)}=1-C_{v,\,h}(x)^{2}\leq 1\,.$}
$\Box$
4.2 Pólya’s theorems for $v\neq 1$ and general weights
We obtain general Pólya’s theorems of the form indicated in the subsection heading.
Theorem 4.5 ($\exists\,U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If $A_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $B_{v,h}(1)_{1}$, $C_{v,h}(1)_{1}$ and $D_{h}(1)_{1}$ exist and $B_{h}(1)_{1}\neq 0$, then
$A_{v,\,h}(1)_{1}\in D_{h}(1)_{1}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,.$
Proof. By item 1 of Proposition 4.2 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$. By Proposition 2.20, $A_{v,h}(1)_{1}=C_{v,h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$. Thus by Proposition 2.22, $\big(C_{v,h}(1)_{1}\big)^{2}=1-\frac{B_{v,h}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We trade the non-vanishing of $B_{h}(1)_{1}$ for the existence of $E_{v,h}(1)_{1}$ and obtain the following variant of the last theorem.
Theorem 4.6 ($\exists\,U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If $A_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $B_{v,h}(1)_{1}$, $C_{v,h}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{v,h}(1)_{1}$ exist, then
$A_{v,\,h}(1)_{1}\in D_{h}(1)_{1}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,$,
where for $B_{h}(1)_{1}=0$ we have $B_{v,h}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{v,h}(1)_{1}=0$.
Proof. If $B_{h}(1)_{1}\neq 0$, we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.2, and Propositions 2.20 and 2.19 that also $B_{v,h}(1)_{1}=0$. Item 3 of Proposition 4.2 and Proposition 2.20 give $C_{v,h}(1)_{1}=0$. Finally, item 1 of Proposition 4.2 and Proposition 2.20 give $A_{v,h}(1)_{1}=0$. $\Box$
Theorem 4.7 ($\exists\,U(1)_{2}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If
$c=\sum_{n\geq 1}|h(B_{n})|<1$
and $B_{v,h}(1)_{2}$ and $D_{h}(1)_{2}$ exist, then the following holds. If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$. If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,